# Inexact Restoration via random models for unconstrained noisy optimization

### v2

@@ Abstract.
The authors are members of the INdAM Research Group GNCS. The research that led to the present paper was partially supported by INDAM-GNCS through Progetti di Ricerca 2024 and by PNRR - Missione 4 Istruzione e Ricerca - Componente C2 Investimento 1.1, Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN) funded by the European Commission under the NextGeneration EU programme, project “Advanced optimization METhods for automated central veIn Sign detection in multiple sclerosis from magneTic resonAnce imaging (AMETISTA)”, code: P2022J9SNP, MUR D.D. financing decree n. 1379 of 1st September 2023 (CUP E53D23017980001), project “Numerical Optimization with Adaptive Accuracy and Applications to Machine Learning”, code: 2022N3ZNAX MUR D.D. financing decree n. 973 of 30th June 2023 (CUP B53D23012670006), and by Partenariato esteso FAIR “Future Artificial Intelligence Research” SPOKE 1 Human-Centered AI. Obiettivo 4, Project “Mathematical and Physical approaches to innovative Machine Learning technologies (MaPLe)”, code: EP_FAIR_002, CUP B93C23001750006.
@@ 1. Introduction
where $\bar{f}(x,y,\omega_{y})$ is an estimate of $f(x)$ with noise level $y\in Y$, $\omega_{y}$ denotes a random variable dependent of $y$ and independent of $x$, and $h:Y\rightarrow\mathbb{R}$ is a nonnegative function such that the lower $h(y)$, the higher the accuracy of the function and gradient estimates is in probability. In the ideal case $h(y)=0$, the function estimate $\bar{f}(x,y,\omega_{y})$ coincides with the exact objective value $f(x)$. We show that our problem setting is viable for well-known optimization problems and then design and analyze a new trust-region algorithm that employs random first-order models for optimization. We borrow ideas from the Inexact Restoration approach and adapt such framework, in order to define an algorithm that is well-defined and theoretically well founded even if the random models are inaccurate at some iterations. We denote our algorithm as irerm (Inexact REstoration with Random Models).
@@ 2. Problem formulation
For any $x\in\mathbb{R}^{n}$ and level of noise $y\in Y$, let $\bar{f}(x,y,\omega_{y})$ and $\nabla_{x}\bar{f}(x,y,\omega_{y})$ be random evaluations of $f(x)$ and $\nabla f(x)$, respectively, where $\omega_{y}$ denotes a random variable from a probability space with probability measure $\mathbb{P}$, dependent of $y$ and independent of $x$, that models the noise on the function and its gradient. Regarding the constraint in (1.2), the function $h:Y\rightarrow\mathbb{R}$ is assumed to be non-negative. If $h(y)=0$, then $\bar{f}(x,y,\omega_{y})$ and $\nabla_{x}\bar{f}(x,y,\omega_{y})$ coincide with $f(x)$ and $\nabla f(x)$, respectively. If $h(y)$ is non-zero, its value is related to the level of accuracy that $\bar{f}(x,y,\omega_{y})$ and $\nabla_{x}\bar{f}(x,y,\omega_{y})$ can achieve in probability with respect to the corresponding exact evaluations. We summarize these assumptions in the following.
@@
    Functions $f(x)$ and $\bar{f}(x,y)$ are bounded from below, i.e., there exists $f_{low}$ such that
@@
    $f(x,y,\omega_{y})$ $\geq f_{low},\quad\forall\ x\in\mathbb{R}^{n},\ \forall\ y\in Y.$
@@
where the symbol $p(y)=\lceil 1/y\rceil\in\mathbb{N}$ represents the number of samples for a given $y\in(0,1]$ and $\omega_{y}$ represents the i.i.d. noise realizations of $\xi$, $i=1,\ldots,p(y)$. In this case, sufficient accuracy in probability is attained in the approximations $\bar{f}(x,y),\nabla_{x}\bar{f}(x,y)$ by choosing $y$ small enough [23, §6] and problem (1.1) can be cast into (1.2), for instance by setting $Y=[0,1]$ and $h(y)=\sqrt{y}$, $y\in Y$.
@@
The estimator $\bar{f}(x,y)$ in (2.8) satisfies the first probabilistic condition in (2.1) if the sample size $p(y)$ satisfies (2.10), see [23, §6]. An analogous result holds for the second probabilistic condition in (2.1), under the assumption that the noise in the gradient computation is unbiased, and the variance of the gradient estimate is upper bounded by $V$. Thus, Assumption 2.1(ii) holds with $\bar{\mu}(\alpha)=\sqrt{(1-\alpha)/V}$, $\bar{\mu}_{up}=\sqrt{1/V}$.
@@ 3. An inexact restoration trust region method with random models
    (3.2) $\theta_{k+1}^{t}=\left\{\begin{array}[]{ll}\theta_{k},&\mbox{if}\ \operatorname{Pred}_{k}(\theta_{k})\geq\theta_{k}\delta_{k}\|g_{k}\|\\ \displaystyle\frac{h(y_{k})-h(\widetilde{y}_{k+1})}{\bar{f}_{k}^{\,\dagger}-\bar{f}_{k}^{*}+h(y_{k})-h(\widetilde{y}_{k+1})},&\mbox{otherwise}.\end{array}\right.$
@@
At Step 2 the function estimate $\bar{f}_{k}^{\dagger}=\bar{f}(x_{k},y_{k+1}^{t},\omega_{y_{k+1}^{t}}^{\dagger})$ and the gradient estimate $g_{k}=\nabla_{x}\bar{f}(x_{k},y_{k},\omega_{y_{k+1}}^{g})$ are computed for noise realizations $\omega_{y_{k+1}^{t}}^{\dagger}$ and $\omega_{y_{k+1}}^{g}$, respectively. Then the linear model $m_{k}(p)=\bar{f}_{k}^{\dagger}+g_{k}^{T}p$ is formed, and finally the search direction $p_{k}$ is computed by minimizing the model over the ball of center zero and radius $\delta_{k}$, equivalently
@@
where $\theta\in(0,1)$ and $\bar{f}^{*}_{k}=\bar{f}(x_{k},y_{k+1}^{t},\omega_{y_{k+1}^{t}}^{*})$ is a further function estimate. More precisely, the parameter $\theta_{k+1}^{t}$ is computed so that the following condition holds
@@
Condition (3.10) is not standard for deterministic Inexact Restoration approaches and deserves some comments. Such a condition has been introduced in order to ensure that the sequence $\{\theta_{k}\}$ is bounded away from zero, allowing us to avoid certain technical assumptions, employed in the previous literature, that may not be satisfied in our settings. More precisely, in seminal deterministic algorithms such as [7, 8, 13], it is proved that the sequence $\{\theta_{k}\}$ is bounded away from zero by enforcing a condition on the noisy function values along the iterations. Such a condition is problem dependent and hard to impose unless the function $\bar{f}(x,y,\omega)$ is Lipschitz continuous with respect to $y$ and $\omega$, which is hardly verifiable or satisfied if the function evaluations are random, as is in our case. Likewise, in the stochastic Inexact Restoration algorithms [4–6] for finite-sum minimization, we proved that $\{\theta_{k}\}$ is bounded away from zero assuming that every subsampled function evaluation is uniformly bounded from below and above on a set containing the generated sequence $\{x_{k}\}$. For the sake of generality, we preferred to abandon such assumption and enforce condition (3.10) instead.
@@ 4.1. On the sequence $\{\theta_{k}\}$
(i) We note that $\theta_{0}>0$, and proceed by induction assuming that $\theta_{k}>0$ for some $k>0$. First, suppose that $h(y_{k})=0$. By (3.1), we have that $h(y_{k+1}^{t})=h(\widetilde{y}_{k+1})=0$, which by Assumption 2.1(i) yields $\bar{f}_{k}^{*}=\bar{f}_{k}^{\,\dagger}=f(x_{k})$. Thus, $\operatorname{Pred}_{k}(\theta_{k})={\theta_{k}}\delta_{k}\|g_{k}\|$, the update rule (3.2) gives $\theta_{k+1}^{t}=\theta_{k}$, and $\theta_{k+1}^{t}$ is positive by induction. Now suppose that $h(y_{k})>0$. If inequality $\operatorname{Pred}_{k}(\theta_{k})\geq\theta_{k}\delta_{k}\|g_{k}\|$ holds, then the update rule (3.2) gives $\theta_{k+1}^{t}=\theta_{k}$, and thus $\theta_{k+1}^{t}>0$ by induction. Otherwise, it must be
@@
By the update rule (3.2), we have $\theta_{k+1}^{t}=\theta_{k+1}^{*}$, which is positive from (4.3).
Since Step 4 assigns either $\theta_{k+1}=\theta_{k+1}^{t}$ or $\theta_{k+1}=\theta_{k}$, we conclude that $\theta_{k+1}$ is positive in both cases. Hence, the sequence $\{\theta_{k}\}$ is positive. Moreover, from (3.2) and (4.3) we have $\theta_{k+1}^{t}\leq\theta_{k}$, and $\{\theta_{k}\}_{k\in\mathbb{N}}$ is nonincreasing by Step 4.
(ii) If inequality $\operatorname{Pred}_{k}(\theta_{k})\geq\theta_{k}\delta_{k}\|g_{k}\|$ holds, the update rule (3.2) gives $\theta_{k+1}^{t}=\theta_{k}$, and thus (4.1) trivially holds. Otherwise, $\theta_{k+1}^{t}=\theta_{k+1}^{*}$, and from (4.2) and (4.3) it follows that $\operatorname{Pred}_{k}(\theta)\geq{\theta}\delta_{k}\|g_{k}\|$ if and only if the value $\theta\leq\theta_{k+1}^{*}$ with $\theta_{k+1}^{*}$ given in (4.3). Thus (4.1) is satisfied.
@@ 4.2. Successful iterations
In this section, we introduce conditions that enforce successful iterations, i.e., conditions which guarantee (3.8)–(3.10). The key issue is the characterization of sufficient accuracy on $f$ and $\nabla f$; in the following definition, we formalize such accuracy requirements and call true an iteration where they hold, false otherwise.
@@
Based on Section 2, we remark that true iterations occur with a certain probability if the level of noise is sufficiently small. In the case of finite-sum minimization, iteration $k$ is true in probability provided that the sample sizes used for computing $\bar{f}_{k}^{*},\,\bar{f}_{k}^{\dagger},\,g_{k},\,\bar{f}_{k}^{p}$ satisfy lower bounds of the form (2.5); in the remaining cases of Section 2, iteration $k$ is true in probability for specified numbers of trials employed in the evaluation of $\bar{f}_{k}^{*},\,\bar{f}_{k}^{\dagger},\,g_{k},\,\bar{f}_{k}^{p}$ and complying with inequalities of the form (2.10). By using Taylor expansion and assuming Lipschitz continuity of $\nabla f$, the second condition in (4.4) and (4.5) imply that model $m_{k}(p)=\bar{f}_{k}^{\dagger}+g_{k}^{T}p$ is fully linear in a ball of center $x_{k}$ and radius $\delta_{k}$, according to the definition of fully linear model given e.g. in [15, Definition 3.1].
@@
Trivially (4.7) implies (3.9). Concerning (3.8), we combine the definition of $\operatorname{Pred}_{k}$ in (3.5), the definition of $\operatorname{Ared}_{k}$ in (3.7), inequality (4.1), condition (3.1) and $\theta_{k+1}^{t}\leq\theta_{k}\leq\theta_{0}$ (see Lemma 4.2(i)), so as to obtain
@@
The occurrence that $k$ is true and Assumption 4.5 give
@@
[…]
@@
Suppose that iteration $k$ is successful. Then, we have
$\operatorname{Ared}_{k}(x_{k+1},\theta_{k+1})=\operatorname{Ared}_{k}(x_{k}+p_{k},\theta_{k+1})\geq\eta_{1}\eta_{2}\theta_{k+1}\delta_{k}^{2}.$
@@
(4.10) $\operatorname{Ared}_{k}(x_{k}+p_{k},\theta_{k+1})$$\geq$ $\eta_{1}{\theta_{k+1}}\delta_{k}\|g_{k}\|$
(4.11) $\geq$ $\eta_{1}\eta_{2}\theta_{k+1}\delta_{k}^{2}.$

### v3

@@ Abstract.
The authors are members of the INdAM Research Group GNCS. Benedetta Morini and Simone Rebegoldi are partially supported by INDAM-GNCS through Progetti di Ricerca 2025 (CUP E53C24001950001). Benedetta Morini and Simone Rebegoldi are partially supported by PNRR - Missione 4 Istruzione e Ricerca - Componente C2 Investimento 1.1, Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN) funded by the European Commission under the NextGeneration EU programme, project “Advanced optimization METhods for automated central veIn Sign detection in multiple sclerosis from magneTic resonAnce imaging (AMETISTA)”, code: P2022J9SNP, MUR $D.$ financing decree n. 1379 of 1st September 2023 (CUP E53D23017980001). Benedetta Morini is partially supported by PNRR - Missione 4 Istruzione e Ricerca - Componente C2 Investimento 1.1, Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN) funded by the European Commission under the NextGeneration EU programme, project “Numerical Optimization with Adaptive Accuracy and Applications to Machine Learning”, code: 2022N3ZNAX, MUR $D.$ financing decree n. 973 of 30th June 2023 (CUP B53D23012670006), and by Partenariato esteso FAIR “Future Artificial Intelligence Research” SPOKE 1 Human-Centered AI. Obiettivo 4, Project “Mathematical and Physical approaches to innovative Machine Learning technologies (MaPLe)”, code: EP_FAIR_002, CUP B93C23001750006.
@@ 1. Introduction
where $\bar{f}(x,y,\omega_{y})$ is an estimate of $f(x)$ with noise level $y\in Y$, $\omega_{y}\in\Omega$ denotes a random variable dependent of $y$ and independent of $x$, and $h:Y\rightarrow\mathbb{R}$ is a nonnegative function such that the lower $h(y)$, the higher the accuracy of the function and gradient estimates is in probability. In the ideal case $h(y)=0$, the function estimate $\bar{f}(x,y,\omega_{y})$ coincides with the exact objective value $f(x)$. We show that our problem setting is viable for well-known optimization problems and then design and analyze a new trust-region algorithm that employs random first-order models for optimization. We borrow ideas from the Inexact Restoration approach and adapt such framework, in order to define an algorithm that is well-defined and theoretically well founded even if the random models are inaccurate at some iterations. We denote our algorithm as irerm (Inexact REstoration with Random Models).
@@ 2. Problem formulation
For any $x\in\mathbb{R}^{n}$ and level of noise $y\in Y$, let $\bar{f}(x,y,\omega_{y})$ and $\nabla_{x}\bar{f}(x,y,\omega_{y})$ be random evaluations of $f(x)$ and $\nabla f(x)$, respectively, where $\omega_{y}\in\Omega$ denotes a random variable from a probability space with probability measure $\mathbb{P}$, dependent of $y$ and independent of $x$, that models the noise on the function and its gradient. Regarding the constraint in (1.2), the function $h:Y\rightarrow\mathbb{R}$ is assumed to be non-negative. If $h(y)=0$, then $\bar{f}(x,y,\omega_{y})$ and $\nabla_{x}\bar{f}(x,y,\omega_{y})$ coincide with $f(x)$ and $\nabla f(x)$, respectively. If $h(y)$ is non-zero, its value is related to the level of accuracy that $\bar{f}(x,y,\omega_{y})$ and $\nabla_{x}\bar{f}(x,y,\omega_{y})$ can achieve in probability with respect to the corresponding exact evaluations. We summarize these assumptions in the following.
@@
    Functions $f(x)$ and $\bar{f}(x,y,\omega_{y})$ are bounded from below, i.e., there exists $f_{low}$ such that
@@
    $\bar{f}(x,y,\omega_{y})$ $\geq f_{low},\quad\forall\ x\in\mathbb{R}^{n},\ \forall\ y\in Y,\ \forall\ \omega_{y}\in\Omega.$
@@
where the symbol $p(y)=\lceil 1/y\rceil\in\mathbb{N}$ represents the number of samples for a given $y\in(0,1]$ and $\omega_{y}$ represents the i.i.d. noise realizations of $\xi$, $i=1,\ldots,p(y)$. In this case, sufficient accuracy in probability is attained in the approximations $\bar{f}(x,y,\omega_{y}),\nabla_{x}\bar{f}(x,y,\omega_{y})$ by choosing $y$ small enough [23, §6] and problem (1.1) can be cast into (1.2), for instance by setting $Y=[0,1]$ and $h(y)=\sqrt{y}$, $y\in Y$.
@@
The estimator $\bar{f}(x,y,\omega_{y})$ in (2.8) satisfies the first probabilistic condition in (2.1) if the sample size $p(y)$ satisfies (2.10), see [23, §6]. An analogous result holds for the second probabilistic condition in (2.1), under the assumption that the noise in the gradient computation is unbiased, and the variance of the gradient estimate is upper bounded by $V$. Thus, Assumption 2.1(ii) holds with $\bar{\mu}(\alpha)=\sqrt{(1-\alpha)/V}$, $\bar{\mu}_{up}=\sqrt{1/V}$.
@@ 3. An inexact restoration trust region method with random models
    (3.2) $\theta_{k+1}^{t}=\left\{\begin{array}[]{ll}\theta_{k},\quad&\mbox{if}\ \operatorname{Pred}_{k}(\theta_{k})\geq\theta_{k}\delta_{k}\|g_{k}\|\\ \displaystyle\frac{h(y_{k})-h(\widetilde{y}_{k+1})}{\bar{f}_{k}^{\,\dagger}-\bar{f}_{k}^{*}+h(y_{k})-h(\widetilde{y}_{k+1})},\quad&\mbox{otherwise}.\end{array}\right$.
@@
At Step 2 the function estimate $\bar{f}_{k}^{\dagger}=\bar{f}(x_{k},y_{k+1}^{t},\omega_{y_{k+1}^{t}}^{\dagger})$ and the gradient estimate $g_{k}=\nabla_{x}\bar{f}(x_{k},y_{k+1}^{g},\omega_{y_{k+1}^{g}})$ are computed for noise realizations $\omega_{y_{k+1}^{t}}^{\dagger}$ and $\omega_{y_{k+1}^{g}}$, respectively. Then the linear model $m_{k}(p)=\bar{f}_{k}^{\dagger}+g_{k}^{T}p$ is formed, and finally the search direction $p_{k}$ is computed by minimizing the model over the ball of center zero and radius $\delta_{k}$, equivalently
@@
where $\theta\in(0,1)$ and $\bar{f}^{*}_{k}=\bar{f}(x_{k},y_{k+1}^{t},\omega_{y_{k+1}^{t}}^{*})$ is a further function estimate computed for the noise realization $\omega_{y_{k+1}^{t}}^{*}$. More precisely, the parameter $\theta_{k+1}^{t}$ is computed so that the following condition holds
@@
Condition (3.10) is not standard for deterministic Inexact Restoration approaches and deserves some comments. Such a condition has been introduced in order to ensure that the sequence $\{\theta_{k}\}$ is bounded away from zero, allowing us to avoid certain technical assumptions, employed in the previous literature, that may not be satisfied in our settings. More precisely, in seminal deterministic algorithms such as [7, 8, 13], it is proved that the sequence $\{\theta_{k}\}$ is bounded away from zero by enforcing a condition on the noisy function values along the iterations. Such a condition is problem dependent and hard to impose unless the function $\bar{f}(x,y,\omega)$ is Lipschitz continuous with respect to $y$ and $\omega$, which is hardly verifiable or satisfied if the function evaluations are random, as is in our case. Likewise, in the stochastic Inexact Restoration algorithms [4–6] for finite-sum minimization, it is proved that $\{\theta_{k}\}$ is bounded away from zero assuming that every subsampled function evaluation is uniformly bounded from below and above on a set containing the generated sequence $\{x_{k}\}$. For the sake of generality, we preferred to abandon such an assumption and enforce condition (3.10) instead.
@@ 4.1. On the sequence $\{\theta_{k}\}$
(i) We note that $\theta_{0}>0$, and proceed by induction assuming that $\theta_{k}>0$ for some $k\geq 0$. First, suppose that $h(y_{k})=0$. By (3.1), we have that $h(y_{k+1}^{t})=h(\widetilde{y}_{k+1})=0$, which by Assumption 2.1(i) yields $\bar{f}_{k}^{*}=\bar{f}_{k}^{\,\dagger}=f(x_{k})$. Thus, $\operatorname{Pred}_{k}(\theta_{k})={\theta_{k}}\delta_{k}\|g_{k}\|$, the update rule (3.2) gives $\theta_{k+1}^{t}=\theta_{k}$, and $\theta_{k+1}^{t}$ is positive by induction. Now suppose that $h(y_{k})>0$. If inequality $\operatorname{Pred}_{k}(\theta_{k})\geq\theta_{k}\delta_{k}\|g_{k}\|$ holds, then the update rule (3.2) gives $\theta_{k+1}^{t}=\theta_{k}$, and thus $\theta_{k+1}^{t}>0$ by induction. Otherwise, it must be
@@
By the update rule (3.2), we have $\theta_{k+1}^{t}=\theta_{k+1}^{*}$, which is positive from (4.4). Since Step 4 assigns either $\theta_{k+1}=\theta_{k+1}^{t}$ or $\theta_{k+1}=\theta_{k}$, we conclude that $\theta_{k+1}$ is positive in both cases. Hence, the sequence $\{\theta_{k}\}$ is positive. Moreover, from (3.2) and (4.4) we have $\theta_{k+1}^{t}\leq\theta_{k}$, and $\{\theta_{k}\}_{k\in\mathbb{N}}$ is nonincreasing by Step 4.
(ii) If inequality $\operatorname{Pred}_{k}(\theta_{k})\geq\theta_{k}\delta_{k}\|g_{k}\|$ holds, the update rule (3.2) gives $\theta_{k+1}^{t}=\theta_{k}$, and thus (4.1) trivially holds. Otherwise, the update rule (3.2) gives $\theta_{k+1}^{t}=\theta_{k+1}^{*}$, which satisfies (4.1). Indeed, imposing $\operatorname{Pred}_{k}(\theta)\geq{\theta}\delta_{k}\|g_{k}\|$ yields
$\theta\left(\bar{f}_{k}^{*}-\bar{f}_{k}^{\,\dagger}-(h(y_{k})-h(\widetilde{y}_{k+1}))\right)\geq-(h(y_{k})-h(\widetilde{y}_{k+1})).$
From (4.3), it follows that $\theta$ must satisfy
$\theta\leq\theta_{k+1}^{*}=\frac{h(y_{k})-h(\widetilde{y}_{k+1})}{\bar{f}_{k}^{\,\dagger}-\bar{f}_{k}^{*}+h(y_{k})-h(\widetilde{y}_{k+1})},$
hence condition (4.1) holds.
@@ 4.2. Successful iterations
In this section, we introduce conditions that enforce successful iterations, i.e., conditions which guarantee (3.8)–(3.10). The key issue is the characterization of sufficient accuracy on $f$ and $\nabla f$. In the following definition, we formalize such accuracy requirements and call true an iteration where they hold, false otherwise.
@@
Based on Section 2, we remark that true iterations occur with a certain probability if the level of noise is sufficiently small. In the case of finite-sum minimization, iteration $k$ is true in probability provided that the sample sizes used for computing $\bar{f}_{k}^{*},\,\bar{f}_{k}^{\dagger},\,g_{k},\,\bar{f}_{k}^{p}$ satisfy lower bounds of the form (2.5); in the remaining cases of Section 2, iteration $k$ is true in probability for specified numbers of trials employed in the evaluation of $\bar{f}_{k}^{*},\,\bar{f}_{k}^{\dagger},\,g_{k},\,\bar{f}_{k}^{p}$ and complying with inequalities of the form (2.10). By using Taylor expansion and assuming Lipschitz continuity of $\nabla f$, the second condition in (4.5) and (4.6) imply that model $m_{k}(p)=\bar{f}_{k}^{\dagger}+g_{k}^{T}p$ is fully linear in a ball of center $x_{k}$ and radius $\delta_{k}$, according to the definition of fully linear model [^1] Suppose $\nabla f$ is Lipschitz continuous and let $\kappa_{ef},\kappa_{eg}$ be some positive constants. A function $m_{k}$ is a $(\kappa_{ef},\kappa_{eg})$-fully linear model on $B=\{x:\|x-x_{k}\|\leq\delta_{k}\}$ if $\forall y\in B$ $\|\nabla f(y)-\nabla m_{k}(y)\|\leq\kappa_{eg}\delta_{k}\ \ \mbox{ and }\ \ |f(y)-m_{k}(y)|\leq\kappa_{ef}\delta_{k}^{2}.$ See e.g., [15, Definition 3.1]. .
@@
Trivially (4.8) implies (3.9). Concerning (3.8), we combine the definition of $\operatorname{Pred}_{k}$ in (3.5), the definition of $\operatorname{Ared}_{k}$ in (3.7), inequality (4.1), condition (3.1) so as to obtain
@@
The occurrence that $k$ is true, Assumption 4.5 and $\theta_{k+1}^{t}\leq\theta_{k}\leq\theta_{0}$ (see Lemma 4.2(i)) give
@@
$+\theta_{0}\left(\|g_{k}-\nabla f(x_{k})\|\delta_{k}+\left|\int_{0}^{1}\left(\nabla f(x_{k})-\nabla f(x_{k}+\tau p_{k})\right)^{T}p_{k}d\tau\right|\right)$
@@
Suppose Assumption 2.1 holds. If the iteration $k$ is successful, then we have
(4.11) $\operatorname{Ared}_{k}(x_{k+1},\theta_{k+1})=\operatorname{Ared}_{k}(x_{k}+p_{k},\theta_{k+1})\geq\eta_{1}\delta_{k}\theta_{k+1}\|g_{k}\|.$
@@
$\operatorname{Ared}_{k}(x_{k}+p_{k},\theta_{k+1})\geq\eta_{1}\operatorname{Pred}_{k}(\theta_{k+1})\geq\eta_{1}{\theta_{k+1}}\delta_{k}\|g_{k}\|.$